Ahmer walks 350m south and 23m west 80m north find the displacement and distance?
1
Expert's answer
2020-03-10T11:28:51-0400

The distance


"d=350+23+80=453\:\rm m."

The displacement


"s=\sqrt{(350-80)^2+23^2}=271\:\rm m."